OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

halp

OpenStudy (anonymous):

like this?\[\frac{ 4 }{ 5x } +\frac{ 4 }{ 3 } = 2x\]

OpenStudy (anonymous):

I guess, it says solve for x

OpenStudy (anonymous):

i know solve for x, but we need to get the equations right first

OpenStudy (anonymous):

oh right lol, sorry im like so out of it when it comes to math Its x=4/5 x + 4/3 = 2x

OpenStudy (anonymous):

the x is separate

OpenStudy (anonymous):

ok lets multiply by 3 on both sides to remove the denominator on the 4/3

OpenStudy (anonymous):

\[(3)\frac{ 4 }{ 5x } + (3) \frac{ 4 }{ 3 } = (3)2x\]

OpenStudy (anonymous):

2.4 x + 4 =6x is that right?

OpenStudy (anonymous):

where did you get the 2.4x?

OpenStudy (anonymous):

3 times 4/5

OpenStudy (anonymous):

its 3 times 4/5x

OpenStudy (anonymous):

12/5x you cant just pull the x out

OpenStudy (anonymous):

i told you that the x was separate though

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }x\]

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

\[2.4x + 4 = 2x\] now minus 2x from both sides

OpenStudy (anonymous):

0.4x + 2x

OpenStudy (anonymous):

or 0.4 + 2x = 2x

OpenStudy (anonymous):

\[2.4x + 4 -2x = 2x - 2x\]

OpenStudy (anonymous):

so would 2x -2x or just 2x be the answer..?

OpenStudy (anonymous):

solve that chunk first

OpenStudy (anonymous):

2x-2x is definitely 0

OpenStudy (anonymous):

XD ye ik i got 4.4?

OpenStudy (anonymous):

2.4x-2x =?

OpenStudy (anonymous):

well that would be 0.4 but i added the 4 cuz u said it up there

OpenStudy (anonymous):

its 2.4x + 4

OpenStudy (anonymous):

not 2.4x +4x

OpenStudy (anonymous):

you can't add the variables and the constants together

OpenStudy (anonymous):

ok then 6.4?

OpenStudy (usukidoll):

"you can't add the variables and the constants together " leave it as 2.4x + 4

OpenStudy (anonymous):

2.4x + 4

OpenStudy (anonymous):

2.4 is a coefficient not a constant term

OpenStudy (anonymous):

so we now have \[0.4x + 4 = 0\]

OpenStudy (anonymous):

so that would be the final answer?

OpenStudy (usukidoll):

ick decimals... I would multiply the whole equation by 10 and then solve for x

OpenStudy (anonymous):

yes multiply by 10

OpenStudy (usukidoll):

decimals are evil. we use multiplication to modify the equation and make life easier.

OpenStudy (anonymous):

hint: calculator doesn't care about decimals

OpenStudy (anonymous):

waddaya got?

OpenStudy (usukidoll):

ghostbusters.

OpenStudy (anonymous):

you seen the goat balls yet?

OpenStudy (usukidoll):

ya ._,

OpenStudy (anonymous):

lel

OpenStudy (anonymous):

Um...so the answer is..0.4x + 4 = 0 also, goat...wtf...

OpenStudy (anonymous):

correct?

OpenStudy (usukidoll):

gotta solve for x =_= that's just an equation

OpenStudy (anonymous):

thats not your final answer

OpenStudy (anonymous):

show it to him if he didn't finish it in 10mins

OpenStudy (usukidoll):

k/

OpenStudy (usukidoll):

so first we multiply the entire equation by 10 because dealing with decimals (especially by hand) is torture 0.4x + 4 = 0 (10)(0.4x)+4(10)=0(10) 4x+40=0 now we need x by itself. We don't want that 40 on the left side, so subtract 40 on both sides finally divide both sides by 4

OpenStudy (anonymous):

\[4x + 40 =0\]\[x + 10 -10 = 0 - 10\]